OpenStudy (anonymous):

(5,1) is a solution to the equation 2x-3y=7 what does this mean?

OpenStudy (anonymous):

it means if you plug in a 5 for x and a 1 for y you get a true statement.

OpenStudy (anonymous):

2(5) - 3(1) = 7 10 - 3 = 7 It's true!

OpenStudy (anonymous):

Consistent even...

OpenStudy (anonymous):

You say tomato. I say tomato.

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

(-3, ) is a solution of y = 4-3x

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

(5,1) is a solution to the equation Doesn't mean it is the only solution....

OpenStudy (anonymous):

The 'huh' wasn't directed to you estudier. It was in response to the second question.

OpenStudy (anonymous):

I was responding to him.... We just crossed.....

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

I assumed he had found another solution....

OpenStudy (anonymous):

i am a she

OpenStudy (anonymous):

If u say so..

OpenStudy (anonymous):

(s)he then

OpenStudy (anonymous):

two different lines are parallel if their slopes are _________ .